An express train travels 150 km in the same time a tourist bus travels 120 km. If the bus travelled 20 kph less than the train, what is the speed of each?
Answer by mananth(16949)

train x 150 miles
Bus x -20 120 miles

t=d/r
150 / x=120/( x-20)
Cross multiply
150 ( x + -20 ) = 120 * x
150 x + -3000 = 120 x
120 x -150 x = -3000
-30 x = -3000
/ -30
x = 100 train mph
80 Bus mph